an angle its complement always have a sum of 90 degrees the sum of the measures of the angle and five times its complement is 298 degrees. What is the measure of the angle?
Answer by asha(30)

let the angle be x and its complement be y
given x+y =90 eqn.1
given x+5y =298 eqn.2
eqn.2-eqn,1 gives
4y =298-90 =208
y =208/4 =52
x=90-52 =38
the angles are 38 degrees and 52 degrees.
good luck !!!